H. Cochran, Inc., is considering a new three-year expansion project that requires an initial fixed asset investment of $2,710,000. The fixed asset will be depreciated straight-line to zero over its three-year tax life, after which time it will be worthless. The project is estimated to generate $3,370,000 in annual sales, with costs of $2,190,000. Assume the tax rate is 24 percent and the required return on the project is 12 percent. What is the project’s NPV?
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(2,710,000/3)=$903333.333

OCF=(Sales-Costs)(1-tax rate)+Tax savings on Annual depreciation

=(3,370,000-2,190,000)(1-0.24)+(0.24*903333.333)

=$1113600

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=1113600[1-(1.12)^-3]/0.12

=1113600*2.40183127

=2674679.3

NPV=Present value of inflows-Present value of outflows

=2674679.3-2,710,000

=($35320.7)(Approx)(Negative).